OpenStudy (anonymous):

show that for all values of m, the line y=mx-3m^2 touches the parabola x^2=12y

OpenStudy (campbell_st):

find the point of intersection by substituting the linear equation into the parabola... and solve for x...

OpenStudy (campbell_st):

the number of solutions will determine if it touches in cuts twice

OpenStudy (campbell_st):

oops or doesn't intersect at all

OpenStudy (anonymous):

u mean sub y into the x square = 12 y ?

OpenStudy (campbell_st):

this is what I mean \[x^2 = 12(mx - 3m^2)\]

OpenStudy (anonymous):

ok coni=tinue

OpenStudy (anonymous):

then do what ?

OpenStudy (anonymous):

show not draw lol , its a calculator free paper dude

OpenStudy (dumbcow):

you have to show that the line then is the tangent line of parabola slope of tangent line = dy/dx dy/dx = x/6 tangent line: Y = (x1/6)(X-x1) + y1 for some point (x1,y1) Y = (x1/6)X + (y1 -x1^2/6) and y1 = x1^2/12 --> Y = (x1/6)X -x1^2/12 = mX -3m^2 m = x/6 for every x value there is a "m" value, thus the line is tangent to parabola for all "m" sorry if this is hard to read

OpenStudy (campbell_st):

so then it becomes \[x^2 = 12mx - 36m^2\] or \[x^2 - 12mx + 36m^2 = 0\] so now solve for x by factoring

OpenStudy (anonymous):

ok wait for a while

OpenStudy (anonymous):

after I got x then what again ?

OpenStudy (campbell_st):

ans when you factor it you get \[(x - 6m)^2 = 0\] solve for x, which seems to show the 2 curves have 1 common point of contact...

OpenStudy (campbell_st):

so you have proven the case...

OpenStudy (anonymous):

ok